using values given whag is value of slope lf I1 vs Vs graph
Assume the following values: R=722 R2-222 Rz=102 R=722 V. = 3 V
2. at junction f what expression is correct? I3=I1+I2 , I2=I1+I3 or I1=I2+I3 ?
Answer #1

Ra=7 R3=lor W Z R2 =21 - Vo = 3V w LH Loit M OV R=7 2. RIER are in series Combination So R14 =R1+Ra=7+7=14r. vs 145 P lor - cRa=7 R3=lor W Z R2 =21 - Vo = 3V w LH Loit M OV R=7 2. RIER are in series Combination So R14 =R1+Ra=7+7=14r. vs 145 P lor - c